# Efficient Feature Aggregation and Scale-Aware Regression for Monocular 3D Object Detection

### v2

@@ III-D Loss Function
1) Selective Confidence-Guided Heatmap Loss: Easy samples, which are typically fully visible and closer to the camera, exhibit less scale complexity. As a result, our scale-aware mechanism struggles to provide additional focus on these targets. To address this issue, we propose the Selective Confidence-Guided Heatmap Loss, which encourages the model to focus more on high-confidence samples, given that the network tends to assign higher confidence scores to easy samples (as shown in Fig. 5). Specifically, this loss function uses an improved version of Focal Loss[62] as the primary loss and introduces a Selective Confidence-Guided (SCG) Loss as an auxiliary loss (Eqn. 15) to enhance training.
@@
The 2D object loss follows the design of CenterNet[18]. It includes the Selective Confidence-Guided Heatmap Loss ${\mathcal{L}_{Heatmap}}$, 2D center offset loss ${\mathcal{L}_{{O_{2D}}}}$ and 2D size loss ${\mathcal{L}_{{S_{2D}}}}$:
@@
Following the MonoDLE paradigm[14], the 3D object loss trains the network to predict key 3D attributes. It includes the 3D center offset loss ${\mathcal{L}_{{O_{3D}}}}$ and 3D size loss ${\mathcal{L}_{{S_{3D}}}}$ with L1 loss, the orientation loss ${\mathcal{L}_{\theta}}$ using multi-bin loss[63], and the depth loss ${\mathcal{L}_{D}}$ supervised by Laplacian aleatoric uncertainty loss[64]:
@@ IV-A Setup
1) Dataset: To evaluate the performance of our proposed method, we conducted experiments on the KITTI 3D object detection benchmark[33] and Waymo Open dataset[34]. The KITTI dataset consists of 7,481 training images and 7,518 testing images. It includes three primary object classes: Car, Pedestrian, and Cyclist, with varying levels of difficulty (Easy, Moderate, and Hard) based on factors such as object size, occlusion, and truncation. Following the previous studies[65], the training images are further split into a training set of 3,712 images and a validation set of 3,769 images.
@@
The Waymo Open Dataset[34] is a pivotal resource for autonomous driving scene understanding, featuring 798 training and 202 validation sequences, yielding approximately 160,000 and 40,000 samples, respectively. For our experiments, we adopt the approach from[65], generating 52,386 training images and 39,848 validation images by sampling every third frame from the training sequences. This dataset captures diverse real-world driving scenarios and categorizes objects into LEVEL_1 and LEVEL_2 based on LiDAR point density.
@@
2) Evaluation metrics: In KITTI, model’s performance is typically measured using average precision in 3D space and birds-eye view ($A{P_{3D}}$ and $A{P_{BEV}}$) at 40 recall positions. The AP is calculated based on Intersection over Union between the predicted 3D bounding boxes and the ground truth, with specific IoU thresholds for each object class – 0.7 for cars and 0.5 for pedestrians and cyclists. Following the official protocol[33], we use $A{P_{3D}}$ and $A{P_{BEV}}$ on Moderate category as main metrics. In Waymo, evaluations use two IoU thresholds (0.5 and 0.7) across four distance ranges: Overall, 0 - 30m, 30 - 50m, and 50m - $\infty$.
@@
3) Data augmentation: To enhance the generalization ability of our model, we incorporated several data augmentation techniques commonly used in object detection. During training, we applied random horizontal flipping and random cropping with a probability of 0.5, random scaling with a probability of 0.4, and random shifting with a probability of 0.1. Additionally, we utilized the MixUp3D technique from MonoLSS[17] as an extra data augmentation strategy to further improve the model’s robustness against occlusions.
@@
4) Implementation details: Our proposed MonoASRH was trained on two RTX 3090 GPUs with batch size of 16. For the EH-FAM, we set the channel dimension ${d_{v}}$ of ${V_{i}}$ to 128 in the multi-head self-attention. Following[21], the RoI-Align size $d\times d$ is set to $7\times 7$. For KITTI, we trained the model for 450 epochs using the Adam optimizer[66], with a weight decay of $51e-$ 5. The initial learning rate was set to $31e-$ 3 and decayed by a factor of 10 at the 250th and 370th epochs. For Waymo Open dataset, we trained the model for only 80 epochs with a batch size of 32, while keeping all other settings consistent with those used for KITTI. During testing, the maximum number of detectable objects per image, $K$, was set to 50.
@@ IV-B Main Results
We compared our proposed MonoASRH with state-of-the-art monocular 3D object detection methods on the KITTI test set for the car category. As shown in Table I, our method outperformed the CAD model-based DCD[10], even without additional data. Compared to other methods that do not use extra data, MonoASRH achieved improvements in $A{P_{3D|R40}}$ across all three difficulty levels by 0.65%, 1.00%, and 1.40%. Experimental results validate the effectiveness of our method in detecting objects in the Car category.
@@
D4LCN[39] CVPR 2020 Depth 16.65 11.72 9.51 22.51 16.02 12.55 -
Kinematic3D[11] ECCV 2020 Multi-frames 19.07 12.72 9.17 26.99 17.52 13.10 120
MonoRUn[3] CVPR 2021 LiDAR 19.65 12.30 10.58 27.94 17.34 15.24 70
CaDDN[4] CVPR 2021 LiDAR 19.17 13.41 11.46 27.94 18.91 17.19 630
AutoShape[9] ICCV 2021 CAD 22.47 14.17 11.36 30.66 20.08 15.59 -
MonoDTR[5] CVPR 2022 LiDAR 21.99 15.39 12.73 28.59 20.38 17.14 37
DCD[10] ECCV 2022 CAD 23.81 15.90 13.21 32.55 21.50 18.25 -
SMOKE[13] CVPRW 2020 None 14.03 9.76 7.84 20.83 14.49 12.75 30
MonoDLE[14] CVPR 2021 None 17.23 12.26 10.29 24.79 18.89 16.00 40
GUPNet[21] ICCV 2021 None 20.11 14.20 11.77 - - - 34
MonoGround[25] CVPR 2022 None 21.37 14.36 12.62 30.07 20.47 17.74 30
DEVIANT[22] ECCV 2022 None 21.88 14.46 11.89 29.65 20.44 17.43 40
MonoCon[16] AAAI 2022 None 22.50 16.46 13.95 31.12 22.10 19.00 26
MonoDDE[24] CVPR 2022 None 24.93 17.14 15.10 33.58 23.46 20.37 40
PDR[67] TCSVT 2023 None 23.69 16.14 13.78 31.76 21.74 18.79 29
DVDET[68] RAL 2023 None 23.19 15.44 13.07 32.05 22.15 19.32 -
MonoDETR[48] ICCV 2023 None 25.00 16.47 13.58 33.60 22.11 18.60 38
MonoLSS[17] 3DV 2024 None 24.91 18.17 15.52 34.90 25.65 22.45 35
MonoCD[27] CVPR 2024 None 25.53 16.59 14.53 33.41 22.81 19.57 36
FD3D[69] AAAI 2024 None 25.38 17.12 14.50 34.20 23.72 20.76 40
@@
CaDDN[4] LiDAR 12.87 8.14 6.76 7.00 3.41 3.30
MonoDTR[5] LiDAR 15.33 10.18 8.61 5.05 3.27 3.19
DCD[10] CAD 10.37 6.73 6.28 4.27 2.74 2.41
DEVIANT[22] None 13.43 8.65 7.69 5.05 3.13 2.59
MonoCon[16] None 13.10 8.41 6.94 2.80 1.92 1.55
MonoDDE[24] None 11.13 7.32 6.67 5.94 3.78 3.33
PDR[67] None 11.61 7.72 6.40 2.72 1.57 1.50
MonoDETR[48] None 12.54 7.89 6.65 7.33 4.18 2.92
@@
GUPNet[21] 22.76 16.46 13.72 31.07 22.94 19.75
DEVIANT[22] 24.63 16.54 14.52 32.60 23.04 19.99
MonoDDE[24] 26.66 19.75 16.72 35.51 26.48 23.07
PDR[67] 27.65 19.44 16.24 35.59 25.72 21.35
MonoDETR[48] 28.84 20.61 16.38 - - -
MonoCD[27] 26.45 19.37 16.38 34.60 24.96 21.51
MonoLSS[17] 25.91 18.29 15.94 34.70 25.36 21.84
FD3D[69] 28.22 20.23 17.04 36.98 26.77 23.16
@@
The results for the Pedestrian and Cyclist categories in the KITTI test set are presented in Table II. For Pedestrians, MonoASRH achieved the best performance across all difficulty levels, outperforming the LiDAR-based MonoDTR method[5]. In the Cyclist category, MonoASRH ranked third at the Easy level but still showed competitive performance, with a slight drop at the Moderate level compared to the Transformer-based MonoDETR[48]. However, it achieved the best performance at the Hard level with a 3.51% $A{P_{3D|R40}}$. These results highlight the challenges of detecting smaller, non-rigid objects like Pedestrians and Cyclists. Nonetheless, MonoASRH’s scale-aware dynamic receptive field adjustment capability enabled strong performance in these categories, validating the model’s generalization and scalability. Additionally, the decoupled convolutional layers in EH-FAM are better equipped to capture the features of these categories with large aspect ratios.
@@
TABLE IV: Comparisons for the Vehicle Category on the Waymo Open validation set. We Use $A{P_{3D}}$ (LEVEL_1 and LEVEL_2, IoU $>$ 0.5 and IoU $>$ 0.7) according to Three Object Distance Intervals. ‘’ denote Reproduced Results.
Methods Reference LEVEL_1(IoU $>$ 0.5) LEVEL_2(IoU $>$ 0.5) LEVEL_1(IoU $>$ 0.7) LEVEL_2(IoU $>$ 0.7)
@@
PatchNet[43] ECCV 2020 2.92 10.03 1.09 0.23 2.42 10.01 1.07 0.22 0.39 1.67 0.13 0.03 0.38 1.67 0.13 0.03
PCT[70] NIPS 2021 4.20 14.70 1.78 0.39 4.03 14.67 1.74 0.36 0.89 3.18 0.27 0.07 0.66 3.18 0.27 0.07
NF-DVT[71] JAS 2024 11.32 23.35 5.76 0.98 11.24 23.20 5.68 0.92 2.76 5.89 0.99 0.11 2.64 5.85 0.96 0.10
@@
GUPNet[21] ICCV 2021 10.02 24.78 4.84 0.22 9.39 24.69 4.67 0.19 2.28 6.15 0.81 0.03 2.14 6.13 0.78 0.02
MonoJSG[72] CVPR 2022 5.65 20.86 3.91 0.97 5.34 20.79 3.79 0.85 0.97 4.65 0.55 0.10 0.91 4.64 0.55 0.09
DEVIANT[22] ECCV 2022 10.98 26.85 5.13 0.18 10.29 26.75 4.95 0.16 2.69 6.95 0.99 0.02 2.52 6.93 0.95 0.02
MonoRCNN++[73] WACV 2023 11.37 27.95 4.07 0.42 10.79 27.88 3.98 0.39 4.28 9.84 0.91 0.09 4.05 9.81 0.89 0.08
MonoUNI[74] NIPS 2023 10.98 26.63 4.04 0.57 10.38 26.57 3.95 0.53 3.20 8.61 0.87 0.13 3.04 8.59 0.85 0.12
MonoCD[27] CVPR 2024 11.62 27.21 5.24 0.83 11.14 27.14 5.12 0.75 3.85 8.43 1.00 0.13 3.50 7.53 0.87 0.11
@@
Table IV compares the $AP_{3D}$ metrics for the “Vehicle” category across various methods on the Waymo Open Val Set. In the 30-50m range, our model achieved state-of-the-art performance. Compared to NF-DVT[71], MonoASRH improved the $AP_{3D}$ by 0.4% and 0.02% at IoU thresholds of 0.5 and 0.7, respectively, within the 30-50m range under the LEVEL_1 setting. Additionally, under the LEVEL_2 setting, the $AP_{3D}$ metrics improved by 0.27% and 0.01% at different IoU thresholds within the 30-50m range. These results further validate the effectiveness of MonoASRH in detecting distant and small objects.
@@ IV-C Ablation Studies
In this section, we evaluate the impact of each component of the proposed framework on the network’s performance. Ablation studies were conducted on the KITTI validation set using the Car category, with $A{P_{3D|R40}}$ and $A{P_{BEV|R40}}$ as metrics. Our baseline model is based on the GUPNet[21] and incorporates the Learnable Sample Selection from MonoLSS[17] for initial training. We compare the effects of different components, including the Adaptive Scale-Aware 3D Regression Head, Efficient Hybrid Feature Aggregation Module, and Selective Confidence-Guided Heatmap Loss. As shown in Fig. 7, we provide detailed visual comparisons of various approaches when the model is not equipped with a specific module. The results in Table V and Fig. 7 demonstrate the contributions of each module to the overall performance of our method.
(a) MonoLSS
(b) MonoDETR
(c) MonoASRH
Figure 9: Qualitative comparison results of 3D object detection on the KITTI val set: (a) MonoLSS[17], (b) MonoDETR[48], and (c) Ours. In each group of images, the first row shows the camera view, and the second row shows the BEV view. In the camera view, predicted and ground truth 3D bounding boxes are shown in purple and green, respectively. In the BEV view, their projections are represented in red and green.
TABLE V: Ablation Study on Different Components of our Overall Framework on KITTI val set for Car Category. ‘$A$’, ‘$E$’ and ‘${\mathcal{L}_{H}}$’ denote Adaptive Scale-Aware 3D Regression Head, Efficient Hybrid Feature Aggregation Module, and Selective Confidence-Guided Heatmap Loss, Respectively.
        $A$        $E$        ${\mathcal{L}_{H}}$        Val, $A{P_{3D|R40}}$(%)
@@
1) Adaptive Scale-Aware 3D Regression Head: In Table VI, we further explore the contribution of each component within the Adaptive Scale-Aware 3D Regression Head to model’s performance. Specifically, we evaluated ASRH under three settings: “w/o Scale-Aware Offset”, where receptive field offsets are derived solely from semantic features without object scale information; “w/o Adaptive Attention Mask”, where the attention mask is randomly initialized; “w/o Attentive Normalization”, where the AN layer[61] is not applied to normalize feature maps ${F_{de}}$. The experimental results show that incorporating scale information is crucial in ASRH, enabling the model to perform dynamic scale-aware adjustments of the receptive field. This significantly improves detection, particularly for small and occluded objects at Moderate and Hard difficulty levels. We also provide detailed visual comparisons in Fig. 8, which further demonstrate the role of each component in ASRH.
@@
Additionally, we compared ASRH with the standard deformable convolution[30]. We inserted deformable convolution into the first layer of each 3D regression head in the baseline model and conducted comparison experiments with ASRH. As shown in Table VII, ASRH consistently outperformed standard deformable convolutions in monocular 3D detection tasks.
TABLE VII: Comparison Experiment between ASRH and Deformable Convolution on KITTI val set for Car Category. ‘$\operatorname{Deform}.$’ Denotes Deformable Convolution.
@@
2) Efficient Hybrid Feature Aggregation Module: In order to validate EH-FAM’s generalizability, we replaced the original feature aggregation layers in three baseline models (MonoDLE[14], GUPNet[21], and MonoLSS[17]) with our EH-FAM:
@@
  • GUPNet employs DLAUp[32], a hierarchical aggregation module. We substituted DLAUp entirely with EH-FAM, preserving the backbone and 3D regression head.

### v3

@@ III-D Loss Function
1) Selective Confidence-Guided Heatmap Loss: Easy samples, which are typically fully visible and closer to the camera, exhibit less scale complexity. As a result, our scale-aware mechanism struggles to provide additional focus on these targets. To address this issue, we propose the Selective Confidence-Guided Heatmap Loss, which encourages the model to focus more on high-confidence samples, given that the network tends to assign higher confidence scores to easy samples (as shown in Fig. 7). Specifically, this loss function uses an improved version of Focal Loss[law2018cornernet] as the primary loss and introduces a Selective Confidence-Guided (SCG) Loss as an auxiliary loss (Eqn. 15) to enhance training.
@@
The 2D object loss follows the design of CenterNet[duan2019centernet]. It includes the Selective Confidence-Guided Heatmap Loss $\mathcal{L}_{Heatmap}$}, 2D center offset loss $\mathcal{L}_{{O_{2D}}}$} and 2D size loss $\mathcal{L}_{{S_{2D}}}$}:
@@
Following the MonoDLE paradigm[ma2021delving], the 3D object loss trains the network to predict key 3D attributes. It includes the 3D center offset loss $\mathcal{L}_{{O_{3D}}}$} and 3D size loss $\mathcal{L}_{{S_{3D}}}$} with L1 loss, the orientation loss $\mathcal{L}_{\theta}$} using multi-bin loss[mousavian20173d], and the depth loss $\mathcal{L}_{D}$} supervised by Laplacian aleatoric uncertainty loss[chen2020monopair]:
@@ IV-A Setup
1) Dataset: To evaluate the performance of our proposed method, we conducted experiments on the KITTI 3D object detection benchmark[geiger2012we] and Waymo Open dataset[sun2020scalability]. The KITTI dataset consists of 7,481 training images and 7,518 testing images. It includes three primary object classes: Car, Pedestrian, and Cyclist, with varying levels of difficulty (Easy, Moderate, and Hard) based on factors such as object size, occlusion, and truncation. Following the previous studies[chen20153d], the training images are further split into a training set of 3,712 images and a validation set of 3,769 images.
@@
The Waymo Open Dataset[sun2020scalability] is a pivotal resource for autonomous driving scene understanding, featuring 798 training and 202 validation sequences, yielding approximately 160,000 and 40,000 samples, respectively. For our experiments, we adopt the approach from[chen20153d], generating 52,386 training images and 39,848 validation images by sampling every third frame from the training sequences. This dataset captures diverse real-world driving scenarios and categorizes objects into LEVEL_1 and LEVEL_2 based on LiDAR point density.
@@
2) Evaluation metrics: In KITTI, model’s performance is typically measured using average precision in 3D space and birds-eye view ($A{P_{3D}}$ and $A{P_{BEV}}$) at 40 recall positions. The AP is calculated based on Intersection over Union between the predicted 3D bounding boxes and the ground truth, with specific IoU thresholds for each object class – 0.7 for cars and 0.5 for pedestrians and cyclists. Following the official protocol[geiger2012we], we use $A{P_{3D}}$ and $A{P_{BEV}}$ on Moderate category as main metrics. In Waymo, evaluations use two IoU thresholds (0.5 and 0.7) across four distance ranges: Overall, 0 - 30m, 30 - 50m, and 50m - $\infty$.
@@
3) Data augmentation: To enhance the generalization ability of our model, we incorporated several data augmentation techniques commonly used in object detection. During training, we applied random horizontal flipping and random cropping with a probability of 0.5, random scaling with a probability of 0.4, and random shifting with a probability of 0.1. Additionally, we utilized the MixUp3D technique from MonoLSS[li2024monolss] as an extra data augmentation strategy to further improve the model’s robustness against occlusions.
@@
4) Implementation details: Our proposed MonoASRH was trained on two RTX 3090 GPUs with batch size of 16. For the EH-FAM, we set the channel dimension $d_{v}$} of $V_{i}$} to 128 in the multi-head self-attention. Following[lu2021geometry], the RoI-Align size $d\times d$ is set to $7\times 7$. For KITTI, we trained the model for 450 epochs using the Adam optimizer[kingma2014adam], with a weight decay of 1e51e-5. The initial learning rate was set to 1e31e-3 and decayed by a factor of 10 at the 250th and 370th epochs. For Waymo Open dataset, we trained the model for only 80 epochs with a batch size of 32, while keeping all other settings consistent with those used for KITTI. During testing, the maximum number of detectable objects per image, $K$, was set to 50.
@@ IV-B Main Results
We compared our proposed MonoASRH with state-of-the-art monocular 3D object detection methods on the KITTI test set for the car category. As shown in Table I, our method outperformed the CAD model-based DCD[li2022densely], even without additional data. Compared to other methods that do not use extra data, MonoASRH achieved improvements in $A{P_{3D|R40}}$ across all three difficulty levels by 0.65%, 1.00%, and 1.40%. Experimental results validate the effectiveness of our method in detecting objects in the Car category.
@@
D4LCN[ding2020learning] CVPR 2020 Depth 16.65 11.72 9.51 22.51 16.02 12.55 -
Kinematic3D[brazil2020kinematic] ECCV 2020 Multi-frames 19.07 12.72 9.17 26.99 17.52 13.10 120
MonoRUn[chen2021monorun] CVPR 2021 LiDAR 19.65 12.30 10.58 27.94 17.34 15.24 70
CaDDN[reading2021categorical] CVPR 2021 LiDAR 19.17 13.41 11.46 27.94 18.91 17.19 630
AutoShape[liu2021autoshape] ICCV 2021 CAD 22.47 14.17 11.36 30.66 20.08 15.59 -
MonoDTR[huang2022monodtr] CVPR 2022 LiDAR 21.99 15.39 12.73 28.59 20.38 17.14 37
DCD[li2022densely] ECCV 2022 CAD 23.81 15.90 13.21 32.55 21.50 18.25 -
SMOKE[liu2020smoke] CVPRW 2020 None 14.03 9.76 7.84 20.83 14.49 12.75 30
MonoDLE[ma2021delving] CVPR 2021 None 17.23 12.26 10.29 24.79 18.89 16.00 40
GUPNet[lu2021geometry] ICCV 2021 None 20.11 14.20 11.77 - - - 34
MonoGround[qin2022monoground] CVPR 2022 None 21.37 14.36 12.62 30.07 20.47 17.74 30
DEVIANT[kumar2022deviant] ECCV 2022 None 21.88 14.46 11.89 29.65 20.44 17.43 40
MonoCon[liu2022learning] AAAI 2022 None 22.50 16.46 13.95 31.12 22.10 19.00 26
MonoDDE[li2022diversity] CVPR 2022 None 24.93 17.14 15.10 33.58 23.46 20.37 40
PDR[sheng2023pdr] TCSVT 2023 None 23.69 16.14 13.78 31.76 21.74 18.79 29
DVDET[hu2023aerial] RAL 2023 None 23.19 15.44 13.07 32.05 22.15 19.32 -
MonoDETR[zhang2023monodetr] ICCV 2023 None 25.00 16.47 13.58 33.60 22.11 18.60 38
MonoLSS[li2024monolss] 3DV 2024 None 24.91 18.17 15.52 34.90 25.65 22.45 35
MonoCD[yan2024monocd] CVPR 2024 None 25.53 16.59 14.53 33.41 22.81 19.57 36
FD3D[wu2024fd3d] AAAI 2024 None 25.38 17.12 14.50 34.20 23.72 20.76 40
@@
CaDDN[reading2021categorical] LiDAR 12.87 8.14 6.76 7.00 3.41 3.30
MonoDTR[huang2022monodtr] LiDAR 15.33 10.18 8.61 5.05 3.27 3.19
DCD[li2022densely] CAD 10.37 6.73 6.28 4.27 2.74 2.41
DEVIANT[kumar2022deviant] None 13.43 8.65 7.69 5.05 3.13 2.59
MonoCon[liu2022learning] None 13.10 8.41 6.94 2.80 1.92 1.55
MonoDDE[li2022diversity] None 11.13 7.32 6.67 5.94 3.78 3.33
PDR[sheng2023pdr] None 11.61 7.72 6.40 2.72 1.57 1.50
MonoDETR[zhang2023monodetr] None 12.54 7.89 6.65 7.33 4.18 2.92
@@
GUPNet[lu2021geometry] 22.76 16.46 13.72 31.07 22.94 19.75
DEVIANT[kumar2022deviant] 24.63 16.54 14.52 32.60 23.04 19.99
MonoDDE[li2022diversity] 26.66 19.75 16.72 35.51 26.48 23.07
PDR[sheng2023pdr] 27.65 19.44 16.24 35.59 25.72 21.35
MonoDETR[zhang2023monodetr] 28.84 20.61 16.38 - - -
MonoCD[yan2024monocd] 26.45 19.37 16.38 34.60 24.96 21.51
MonoLSS[li2024monolss] 25.91 18.29 15.94 34.70 25.36 21.84
FD3D[wu2024fd3d] 28.22 20.23 17.04 36.98 26.77 23.16
@@
The results for the Pedestrian and Cyclist categories in the KITTI test set are presented in Table II. For Pedestrians, MonoASRH achieved the best performance across all difficulty levels, outperforming the LiDAR-based MonoDTR method[huang2022monodtr]. In the Cyclist category, MonoASRH ranked third at the Easy level but still showed competitive performance, with a slight drop at the Moderate level compared to the Transformer-based MonoDETR[zhang2023monodetr]. However, it achieved the best performance at the Hard level with a 3.51% $A{P_{3D|R40}}$. These results highlight the challenges of detecting smaller, non-rigid objects like Pedestrians and Cyclists. Nonetheless, MonoASRH’s scale-aware dynamic receptive field adjustment capability enabled strong performance in these categories, validating the model’s generalization and scalability. Additionally, the decoupled convolutional layers in EH-FAM are better equipped to capture the features of these categories with large aspect ratios.
@@
TABLE IV: Comparisons for the Vehicle Category on the Waymo Open validation set. We Use $A{P_{3D}}$ (LEVEL_1 and LEVEL_2, IoU >> 0.5 and IoU >> 0.7) according to Three Object Distance Intervals. ‘’ denote Reproduced Results.
Methods Reference LEVEL_1(IoU >> 0.5) LEVEL_2(IoU >> 0.5) LEVEL_1(IoU >> 0.7) LEVEL_2(IoU >> 0.7)
@@
PatchNet[ma2020rethinking] ECCV 2020 2.92 10.03 1.09 0.23 2.42 10.01 1.07 0.22 0.39 1.67 0.13 0.03 0.38 1.67 0.13 0.03
PCT[wang2021progressive] NIPS 2021 4.20 14.70 1.78 0.39 4.03 14.67 1.74 0.36 0.89 3.18 0.27 0.07 0.66 3.18 0.27 0.07
NF-DVT[pan2024depth] JAS 2024 11.32 23.35 5.76 0.98 11.24 23.20 5.68 0.92 2.76 5.89 0.99 0.11 2.64 5.85 0.96 0.10
@@
GUPNet[lu2021geometry] ICCV 2021 10.02 24.78 4.84 0.22 9.39 24.69 4.67 0.19 2.28 6.15 0.81 0.03 2.14 6.13 0.78 0.02
MonoJSG[lian2022monojsg] CVPR 2022 5.65 20.86 3.91 0.97 5.34 20.79 3.79 0.85 0.97 4.65 0.55 0.10 0.91 4.64 0.55 0.09
DEVIANT[kumar2022deviant] ECCV 2022 10.98 26.85 5.13 0.18 10.29 26.75 4.95 0.16 2.69 6.95 0.99 0.02 2.52 6.93 0.95 0.02
MonoRCNN++[shi2023multivariate] WACV 2023 11.37 27.95 4.07 0.42 10.79 27.88 3.98 0.39 4.28 9.84 0.91 0.09 4.05 9.81 0.89 0.08
MonoUNI[jinrang2023monouni] NIPS 2023 10.98 26.63 4.04 0.57 10.38 26.57 3.95 0.53 3.20 8.61 0.87 0.13 3.04 8.59 0.85 0.12
MonoCD[yan2024monocd] CVPR 2024 11.62 27.21 5.24 0.83 11.14 27.14 5.12 0.75 3.85 8.43 1.00 0.13 3.50 7.53 0.87 0.11
@@
Table IV compares the $AP_{3D}$ metrics for the “Vehicle” category across various methods on the Waymo Open Val Set. In the 30-50m range, our model achieved state-of-the-art performance. Compared to NF-DVT[pan2024depth], MonoASRH improved the $AP_{3D}$ by 0.4% and 0.02% at IoU thresholds of 0.5 and 0.7, respectively, within the 30-50m range under the LEVEL_1 setting. Additionally, under the LEVEL_2 setting, the $AP_{3D}$ metrics improved by 0.27% and 0.01% at different IoU thresholds within the 30-50m range. These results further validate the effectiveness of MonoASRH in detecting distant and small objects.
@@ IV-C Ablation Studies
In this section, we evaluate the impact of each component of the proposed framework on the network’s performance. Ablation studies were conducted on the KITTI validation set using the Car category, with $A{P_{3D|R40}}$ and $A{P_{BEV|R40}}$ as metrics. Our baseline model is based on the GUPNet[lu2021geometry] and incorporates the Learnable Sample Selection from MonoLSS[li2024monolss] for initial training. We compare the effects of different components, including the Adaptive Scale-Aware 3D Regression Head, Efficient Hybrid Feature Aggregation Module, and Selective Confidence-Guided Heatmap Loss. As shown in Fig. 9, we provide detailed visual comparisons of various approaches when the model is not equipped with a specific module. The results in Table V and Fig. 9 demonstrate the contributions of each module to the overall performance of our method.
TABLE V: Ablation Study on Different Components of our Overall Framework on KITTI val set for Car Category. ‘AA’, ‘$E$’ and $\mathcal{L}_{H}$}’ denote Adaptive Scale-Aware 3D Regression Head, Efficient Hybrid Feature Aggregation Module, and Selective Confidence-Guided Heatmap Loss, Respectively.
    AA    $E$    $\mathcal{L}_{H}$}    Val, $A{P_{3D|R40}}$(%)
@@
1) Adaptive Scale-Aware 3D Regression Head: In Table VI, we further explore the contribution of each component within the Adaptive Scale-Aware 3D Regression Head to model’s performance. Specifically, we evaluated ASRH under three settings: “w/o Scale-Aware Offset”, where receptive field offsets are derived solely from semantic features without object scale information; “w/o Adaptive Attention Mask”, where the attention mask is randomly initialized; “w/o Attentive Normalization”, where the AN layer[li2020attentive] is not applied to normalize feature maps $F_{de}$}. The experimental results show that incorporating scale information is crucial in ASRH, enabling the model to perform dynamic scale-aware adjustments of the receptive field. This significantly improves detection, particularly for small and occluded objects at Moderate and Hard difficulty levels. We also provide detailed visual comparisons in Fig. 10, which further demonstrate the role of each component in ASRH.
@@
Additionally, we compared ASRH with the standard deformable convolution[zhu2019deformable]. We inserted deformable convolution into the first layer of each 3D regression head in the baseline model and conducted comparison experiments with ASRH. As shown in Table VII, ASRH consistently outperformed standard deformable convolutions in monocular 3D detection tasks.
TABLE VII: Comparison Experiment between ASRH and Deformable Convolution on KITTI val set for Car Category. $\operatorname{Deform}.$’ Denotes Deformable Convolution.
@@
2) Efficient Hybrid Feature Aggregation Module: In order to validate EH-FAM’s generalizability, we replaced the original feature aggregation layers in three baseline models (MonoDLE[ma2021delving], GUPNet[lu2021geometry], and MonoLSS[li2024monolss]) with our EH-FAM:
@@
  • GUPNet employs DLAUp[yu2018deep], a hierarchical aggregation module. We substituted DLAUp entirely with EH-FAM, preserving the backbone and 3D regression head.
@@
(a) MonoLSS
(b) MonoDETR
(c) MonoASRH
Figure 11: Qualitative comparison results of 3D object detection on the KITTI val set: (a) MonoLSS[li2024monolss], (b) MonoDETR[zhang2023monodetr], and (c) Ours. In each group of images, the first row shows the camera view, and the second row shows the BEV view. In the camera view, predicted and ground truth 3D bounding boxes are shown in purple and green, respectively. In the BEV view, their projections are represented in red and green.